# LLMs instead of Human Judges? A Large Scale Empirical Study across 20 NLP Evaluation Tasks

### v2

@@ Appendix A Datasets
This section provides brief descriptions of the datasets employed in our study. Table 2 summarises relevant dataset information. Note that dataset sizes as reported in Table 2 refer to the number of annotated samples (not to the total number of collected annotations) and might therefore differ from the figures reported in the original papers.
@@
Dataset Task Size Type Guidelines Expert Agreement Leaked
CoLA Warstadt et al. (2019) Acceptability 1,043 Categorical
CoLA-grammar Warstadt and Bowman (2020) Acceptability 1,043 Categorical
Switchboard Wallbridge et al. (2022) Acceptability 100 Graded
Dailydialog Wallbridge et al. (2022) Acceptability 100 Graded
Inferential strategies Mondorf and Plank (2024) Reasoning 300 Categorical
ROSCOE (Golovneva et al., 2023) Reasoning 756 Categorical + Graded
Recipe-generation (Stein et al., 2023) Planning 52 Graded
Medical-safety (Abercrombie and Rieser, 2022) Toxicity & Safety 3,701 Preference
DICES Aroyo et al. (2023) Toxicity & Safety 1,340 Categorical Mixed
ToxicChat Lin et al. (2023) Toxicity & Safety 5,654 Categorical
Topical Chat Mehri and Eskenazi (2020) Dialogue 60 Graded + Categorical
Persona Chat Mehri and Eskenazi (2020) Dialogue 60 Graded + Categorical
WMT 2020 En-De (Freitag et al., 2021) Machine Translation 14,122 Graded
WMT 2020 Zh-En (Freitag et al., 2021) Machine Translation 19,974 Graded
WMT 2023 En-De (Kocmi et al., 2023) Machine Translation 6,588 Graded
WMT 2023 Zh-En (Kocmi et al., 2023) Machine Translation 13,245 Graded
G-Eval / SummEval (Liu et al., 2023) Summarisation 1,600 Graded
QAGS (Wang et al., 2020) Summarisation 953 Categorical
NewsRoom (Grusky et al., 2018) Summarisation 420 Graded
LLMBar (Zeng et al., 2024) Instruction Following 419 Categorical
Table 2: Overview of the main features of the datasets considered in the study. Note that ‘Size’ refers to the number of annotated samples, not to the total number of human annotations. ‘Agreement’ indicates whether multiple annotations are available for the same instance or not. Information on possible data leakage was retrieved from Balloccu et al. (2024).
@@
These datasets contain human judgments on the quality of machine- and human-generated responses based on the provided dialogue context. The annotated dialogues were selected from Topical Chat (Gopalakrishnan et al., 2019)—a dataset collecting human-human conversations on provided facts—and Persona Chat (Zhang et al., 2018), which contains human-human persona-conditioned conversations. Each response is evaluated on 6 attributes: Understandable, Natural, Maintains Context, Interesting, Uses Knowledge, and Overall Quality.
@@
Appendix B Upper Bound Estimation for Model Correlations
Whenever multiple human annotations were publicly available for a property, we computed upper-bound estimates for the correlations achievable by models. The intuition behind these estimates, borrowed from neuroscience Nili et al. (2014), is that the maximum correlation a model can achieve with aggregated human responses is bounded by the average correlation between single-participant responses and the aggregated responses across participants. We applied a similar logic to the human judgments used in the present study and combined it with a bootstrapping approach. For each annotated property, we bootstrapped single-participant responses by sampling $1000$ times from the available human responses, excluding data points where a single annotation was available. Next, we computed the alignment between each of the bootstrapped-participant arrays and the array of aggregated responses. Alignment was computed as Spearman’s correlation for graded judgments and Cohen’s kappa for categorical judgments. Finally, we estimated the upper bound as the average of the $1000$ alignment measures. In cases where alignment between bootstrapped and aggregated responses could not be computed—because the variance of the bootstrapped responses was null—values were replaced with an average of the ‘non-nan’ correlations.
@@ Appendix B Upper Bound Estimation for Model Correlations
Appendix C Inference Details
@@
Appendix D Valid Response Rates
Table 4 reports the rate of valid responses for each model and dataset. Valid response rates are summarised per model and dataset in Figures 4 and 5.
@@ Appendix D Valid Response Rates
Appendix E More Details on Toxicity and Safety Evaluation
@@
Appendix F Additional Results
In Table 5 we report human-model alignment scores per dataset for all models tested, thus complementing Table 1 in the paper. Figure 6 shows alignment scores broken down according to the source of the material to be judged, i.e., human or machine-generated output.
@@ Appendix F Additional Results
Graded Annotations Dailydialog (1) 0.69 0.6 0.55 0.63 0.63 0.52 0.23 0.61 0.48 0.09 0.07
Switchboard (1) 0.66 0.45 0.63 0.59 0.56 0.36 0.53 0.28 0.52 0.13 0.3
Persona Chat (4) 0.22 ±0.11 -0.02 ±0.2 0.16 ±0.1 0.1 ±0.09 0.02 ±0.15 0.07 ±0.13 0.05 ±0.2 -0.02 ±0.14 -0.09 ±0.17 0.03 ±0.13 -0.06 ±0.14
Topical Chat (4) 0.26 ±0.03 0.28 ±0.1 0.13 ±0.04 0.17 ±0.12 0.21 ±0.18 0.14 ±0.05 0.07 ±0.07 0.15 ±0.13 0.29 ±0.11 0.14 ±0.16 0.08 ±0.21
Recipe-generation (6) 0.78 ±0.05 0.66 ±0.07 0.6 ±0.15 0.67 ±0.09 0.57 ±0.24 0.32 ±0.28 0.06 ±0.26 0.34 ±0.09 0.28 ±0.08 0.04 ±0.17 0.1 ±0.08
ROSCOE-GSM8K (2) 0.82 ±0.12 0.83 ±0.11 0.81 ±0.14 0.81 ±0.12 0.79 ±0.13 0.68 ±0.2 0.7 ±0.08 0.76 ±0.15 0.63 ±0.18 0.46 ±0.13 0.1 ±0.07
ROSCOE-eSNLI (2) 0.49 ±0.24 0.4 ±0.16 0.38 ±0.17 0.35 ±0.21 0.32 ±0.12 0.09 ±0.08 0.28 ±0.21 0.19 ±0.16 0.32 ±0.12 0.11 ±0.06 0.11 ±0.17
ROSCOE-DROP (2) 0.57 ±0.22 0.59 ±0.16 0.44 ±0.15 0.44 ±0.13 0.32 ±0.12 0.21 ±0.22 0.37 ±0.18 0.23 ±0.1 0.22 ±0.22 0.16 ±0.17 0.15 ±0.21
ROSCOE-CosmosQA (2) 0.57 ±0.18 0.55 ±0.18 0.51 ±0.16 0.57 ±0.17 0.53 ±0.21 0.33 ±0.25 0.48 ±0.17 0.44 ±0.26 0.57 ±0.2 0.13 ±0.04 0.49 ±0.24
NewsRoom (4) 0.59 ±0.02 0.59 ±0.03 0.44 ±0.05 0.55 ±0.03 0.5 ±0.07 0.36 ±0.06 0.16 ±0.05 0.45 ±0.04 0.26 ±0.06 0.21 ±0.08 -0.01 ±0.04
SummEval (4) 0.35 ±0.06 0.44 ±0.14 0.54 ±0.08 0.38 ±0.02 0.48 ±0.02 0.19 ±0.06 0.13 ±0.06 0.29 ±0.09 0.4 ±0.12 0.15 ±0.05 0.06 ±0.02
WMT 2020 En-De (1) 0.63 0.37 0.51 0.46 0.2 0.42 0.15 0.11 0.36 0.15 -0.03
WMT 2020 Zh-En (1) 0.54 0.39 0.48 0.41 0.25 0.42 0.15 0.14 0.39 0.15 0.01
WMT 2023 En-De (1) 0.22 0.14 0.23 0.16 0.17 0.22 0.19 0.08 0.18 -0.09 -0.05
WMT 2023 Zh-En (1) 0.17 0.14 0.19 0.14 0.15 0.15 0.14 0.02 0.15 0.01 0.01
Table 5: Scores per dataset for all models we evaluate: Cohen’s kappa for categorical annotations and Spearman’s correlation for graded annotations. Datasets in blue concern human-generated language while those in red concern model-generated text.
@@
Graded Annotations Dailydialog (1) 0.69 0.62 0.56 0.25 0.42 0.51 0.4 0.34 0.27
Switchboard (1) 0.6 0.48 0.53 0.07 0.38 0.17 0.36 0.35 0.08
Persona Chat (4) 0.2 ±0.09 0.09 ±0.2 0.11 ±0.06 0.06 ±0.15 0.17 ±0.21 -0.04 ±0.22 0.04 ±0.13 -0.01 ±0.22 0.06 ±0.18
Topical Chat (4) 0.22 ±0.02 0.14 ±0.13 0.11 ±0.1 0.06 ±0.18 0.1 ±0.14 0.16 ±0.17 0.25 ±0.05 0.17 ±0.09 0.08 ±0.13
Recipe-generation (6) 0.67 ±0.12 0.64 ±0.14 0.65 ±0.09 0.42 ±0.18 0.14 ±0.15 0.31 ±0.15 0.41 ±0.07 0.34 ±0.2 0.09 ±0.1
ROSCOE-GSM8K (2) 0.82 ±0.12 0.81 ±0.12 0.81 ±0.11 0.81 ±0.13 0.49 ±0.13 0.8 ±0.11 0.58 ±0.13 0.64 ±0.15 0.07 ±0.07
ROSCOE-eSNLI (2) 0.49 ±0.29 0.39 ±0.38 0.31 ±0.32 0.31 ±0.09 0.33 ±0.01 0.23 ±0.17 0.17 ±0.03 0.19 ±0.13 -0.05 ±0.07
ROSCOE-DROP (2) 0.54 ±0.17 0.55 ±0.19 0.44 ±0.12 0.29 ±0.14 0.29 ±0.24 0.44 ±0.07 0.28 ±0.03 0.17 ±0.15 -0.04 ±0.02
ROSCOE-CosmosQA (2) 0.57 ±0.21 0.55 ±0.22 0.56 ±0.11 0.55 ±0.12 0.62 ±0.15 0.38 ±0.06 0.54 ±0.21 0.32 ±0.14 0.3 ±0.14
NewsRoom (4) 0.57 ±0.05 0.53 ±0.03 0.53 ±0.05 0.46 ±0.02 0.19 ±0.06 0.49 ±0.04 0.19 ±0.04 0.22 ±0.12 0.09 ±0.06
SummEval (4) 0.45 ±0.11 0.48 ±0.16 0.33 ±0.03 0.35 ±0.05 0.17 ±0.05 0.24 ±0.13 0.38 ±0.1 0.24 ±0.11 0.03 ±0.07
WMT 2020 En-De (1) 0.57 0.44 0.38 0.39 0.13 0.34 0.33 0.3 0.0
WMT 2020 Zh-En (1) 0.52 0.44 0.44 0.42 0.19 0.36 0.39 0.35 0.13
WMT 2023 En-De (1) 0.2 0.16 0.18 0.2 0.2 0.18 0.21 0.19 -0.01
WMT 2023 Zh-En (1) 0.15 0.13 0.15 0.16 0.16 0.13 0.11 0.13 0.06
Table 6: Scores per dataset for all models we evaluate using CoT prompts: Cohen’s kappa for categorical annotations and Spearman’s correlation for graded annotations. Datasets in blue concern human-generated language while those in red concern model-generated text.
Prompt Llama 3.1 8B Llama 3.1 70B Mixtral-8x7B
DICES-350-expert Original 0.01 -0.13 -0.11
CoT -0.07 -0.26 -0.02
Few-shot 0.01 -0.22 -0.01
Paraphrase -0.13 -0.36 -0.09
WMT 2023 En-De Original 0.08 0.14 0.17
CoT 0.34 0.16 0.20
Few-shot 0.19 0.21 0.20
Paraphrase 0.02 ±0.08 0.08 ± 0.12 0.14 ±0.05
WMT 2023 Zh-En Original 0.02 0.14 0.15
CoT 0.36 0.16 0.13
Few-shot 0.15 0.21 0.14
Paraphrase 0.08 ±0.04 0.09 ±0.06 0.13 ±0.03
Table 7: Spearman’s correlation for three datasets with graded annotations, comparing the original prompt and CoT prompt to few-shot prompts and prompt paraphrases for a selection of models. For datasets with more than one paraphrased prompt, we report the average and standard deviation across paraphrases.

### v3

@@ Appendix A Datasets
This section provides brief descriptions of the datasets employed in our study. Table 2 summarises relevant dataset information. Note that dataset sizes as reported in Table 2 refer to the number of annotated samples (not to the total number of collected annotations) and might therefore differ from the figures reported in the original papers. Table 3 reports Krippendorf’s $\alpha$ for those datasets with multiple public human annotations.
@@
Dataset Task Size # Annot. Type Guidelines Expert Leaked
CoLA Warstadt et al. (2019) Acceptability 1,043 - Categorical
CoLA-grammar Warstadt and Bowman (2020) Acceptability 1,043 - Categorical
Switchboard Wallbridge et al. (2022) Acceptability 100 3-6 Graded
Dailydialog Wallbridge et al. (2022) Acceptability 100 3-6 Graded
Inferential strategies Mondorf and Plank (2024) Reasoning 300 - Categorical
ROSCOE (Golovneva et al., 2023) Reasoning 756 - Categorical + Graded
Recipe-generation (Stein et al., 2023) Planning 52 - Graded
Medical-safety (Abercrombie and Rieser, 2022) Toxicity & Safety 3,701 - Preference
DICES Aroyo et al. (2023) Toxicity & Safety 1,340 ~70 + ~120 Categorical Mixed
ToxicChat Lin et al. (2023) Toxicity & Safety 5,654 - Categorical
Topical Chat Mehri and Eskenazi (2020) Dialogue 60 3 Graded + Categorical
Persona Chat Mehri and Eskenazi (2020) Dialogue 60 3 Graded + Categorical
WMT 2020 En-De (Freitag et al., 2021) Machine Translation 14,122 3 Graded
WMT 2020 Zh-En (Freitag et al., 2021) Machine Translation 19,974 3 Graded
WMT 2023 En-De (Kocmi et al., 2023) Machine Translation 6,588 - Graded
WMT 2023 Zh-En (Kocmi et al., 2023) Machine Translation 13,245 - Graded
G-Eval / SummEval (Liu et al., 2023) Summarisation 1,600 - Graded
QAGS (Wang et al., 2020) Summarisation 953 3 Categorical
NewsRoom (Grusky et al., 2018) Summarisation 420 3 Graded
LLMBar (Zeng et al., 2024) Instruction Following 419 - Categorical
Table 2: Overview of the main features of the datasets considered in the study. Note that ‘Size’ refers to the number of annotated samples, not to the total number of human annotations. ‘# Annot.’ refers to the number of available individual annotations, if any, which we use to estimate the human upper bound. Note that datasets with only a single annotation per sample, or which only report the average over multiple annotations are not included in ‘# Annot.’. Information on possible data leakage was retrieved from Balloccu et al. (2024).
@@
These datasets contain human judgments on the quality of machine- and human-generated responses based on the provided dialogue context. The annotated dialogues were selected from Topical Chat (Gopalakrishnan et al., 2019)—a dataset collecting human-human conversations on provided facts—and Persona Chat (Zhang et al., 2018), which contains human-human persona-conditioned conversations. Each response is evaluated on 6 attributes: Understandable, Natural, Maintains Context, Interesting/Engaging, Uses Knowledge, and Overall Quality.
@@
Appendix B The Judge-Bench Data Schema
To facilitate extending our benchmark, we adopt a shared schema used to pre-process all datasets. Our publicly available code base includes an example [^2] https://github.com/dmg-illc/JUDGE-BENCH/blob/master/data/example.json of this format as well as instructions on how to verify that newly added datasets comply with it.
The Json-based Judge-Bench data schema ensures that the following fields are included for each dataset:
  • dataset: the name of the dataset;
  • dataset_url: the URL where the original dataset can be downloaded, as opensourced by their creators;
  • annotations: an overview of the properties annotated for each dataset, along with information on how they are measured and prompt-like instructions similar to those originally given to the human annotators (when applicable);
  • instances: dataset instances including the piece of text to be judged, aggregated human judgments and, when available, individual human annotations.
We note that, while we do not systematically explore inter-annotator variations at the instance level, the data schema we adopt allows for conducting this type of analysis in future work.
Appendix C Upper Bound Estimation for Model Correlations
Whenever multiple human annotations were publicly available for a property (see Table 3 for inter-annotator agreement scores), we computed upper-bound estimates for the correlations achievable by models. The intuition behind these estimates, borrowed from neuroscience Nili et al. (2014), is that the maximum correlation a model can achieve with aggregated human responses is bounded by the average correlation between single-participant responses and the aggregated responses across participants. We applied a similar logic to the human judgments used in the present study and combined it with a bootstrapping approach. For each annotated property, we bootstrapped single-participant responses by sampling $1000$ times from the available human responses, excluding data points where a single annotation was available. Next, we computed the alignment between each of the bootstrapped-participant arrays and the array of aggregated responses. Alignment was computed as Spearman’s correlation for graded judgments and Cohen’s kappa for categorical judgments. Finally, we estimated the upper bound as the average of the $1000$ alignment measures. In cases where alignment between bootstrapped and aggregated responses could not be computed—because the variance of the bootstrapped responses was null—values were replaced with an average of the ‘non-nan’ correlations.
@@ Appendix C Upper Bound Estimation for Model Correlations
Appendix D Properties with Graded Judgments
In Figure 3, we display results for a set of graded properties annotated in one or more of the datasets we consider. The properties are defined as follows:
  • Acceptability refers to whether it is plausible or not that a response belongs to a telephonic dialogue and was annotated in Swithboard and Dailydialog;
  • Coherence was annotated for summaries and model-generated reasonings as part of the datasets NewsRoom, ROSCOE, and SummEval;
  • Consistency refers to the alignment between facts described in a summary and in its source text, and was annotated in SummEval;
  • Engaging indicates whether a response generated in the context of a dialogue is dull or interesting and was annotated in TopicalChat and PersonaChat;
  • Fluency measures whether a piece of text is grammatically correct and well-formatted, and was annotated in NewsRoom, SummEval and Recipe-generation;
  • Informativeness refers to the extent to which a summary captures the key points of the full text, and was annotated for summaries as part of the NewsRoom dataset;
  • Relevance refers to whether a summary selects important information as opposed to including redundancies, and was annotated for NewsRoom and SummEval;
  • Verbosity indicates whether a generated recipe is concise and avoids unnecessary repetitions, and was annotated in Recipe-generation.
Appendix E Inference Details
@@
Appendix F Valid Response Rates
Table 5 reports the rate of valid responses for each model and dataset. Valid response rates are summarised per model and dataset in Figures 5 and 6.
@@ Appendix F Valid Response Rates
Appendix G More Details on Toxicity and Safety Evaluation
@@
Appendix H Additional Results
In Table 6 we report human-model alignment scores per dataset for all models tested, thus complementing Table 1 in the paper.
@@ Appendix H Additional Results
Prompt Llama 3.1 8B Llama 3.1 70B Mixtral-8x7B
DICES-350-expert Original 0.01 -0.13 -0.11
CoT -0.07 -0.26 -0.02
Few-shot 0.01 -0.22 -0.01
Paraphrase -0.13 -0.36 -0.09
WMT 2023 En-De Original 0.08 (1) 0.14 (1) 0.17 (1)
CoT 0.18 (1) 0.16 (1) 0.20 (1)
Few-shot 0.19 (1) 0.20 (1) 0.20 (1)
Paraphrase 0.01 ±0.09 (3) 0.08 ± 0.12 (3) 0.14 ±0.05 (3)
WMT 2023 Zh-En Original 0.02 (1) 0.14 (1) 0.15 (1)
CoT 0.13 (1) 0.13 (1) 0.16 (1)
Few-shot 0.15 (1) 0.14 (1) 0.16 (1)
Paraphrase 0.08 ±0.04 (3) 0.09 ±0.07 (2) 0.13 ±0.03 (3)
Table 4: Cohen’s kappa for DICES-350-expert and Spearman’s correlation for two WMT 2023 datasets, comparing the original prompt and CoT prompt to few-shot prompts and prompt paraphrases for a selection of models. For datasets with more than one paraphrased prompt, we report the average and standard deviation across paraphrases. For Spearman’s correlations, we report the number of significant correlations ($p<0.05$) for each model and dataset in brackets.
@@
Graded Annotations Dailydialog (1) 0.69 (1) 0.6 (1) 0.55 (1) 0.63 (1) 0.63 (1) 0.52 (1) 0.23 (1) 0.61 (1) 0.48 (1) 0.09 (0) 0.07 (0)
Switchboard (1) 0.66 (1) 0.45 (1) 0.63 (1) 0.59 (1) 0.56 (1) 0.36 (1) 0.53 (1) 0.28 (1) 0.52 (1) 0.13 (0) 0.3 (1)
Persona Chat (4) 0.22 ±0.11 (2) -0.02 ±0.2 (0) 0.16 ±0.1 (1) 0.1 ±0.09 (0) 0.02 ±0.15 (0) 0.07 ±0.13 (0) 0.05 ±0.2 (0) -0.02 ±0.14 (0) -0.09 ±0.17 (1) 0.03 ±0.13 (0) -0.06 ±0.14 (0)
Topical Chat (4) 0.26 ±0.03 (2) 0.28 ±0.1 (2) 0.13 ±0.04 (0) 0.17 ±0.12 (1) 0.21 ±0.18 (1) 0.14 ±0.05 (0) 0.07 ±0.07 (0) 0.15 ±0.13 (0) 0.29 ±0.11 (3) 0.14 ±0.16 (1) 0.08 ±0.21 (1)
Recipe-generation (6) 0.78 ±0.05 (6) 0.66 ±0.07 (6) 0.6 ±0.15 (6) 0.67 ±0.09 (5) 0.57 ±0.24 (5) 0.32 ±0.28 (5) 0.06 ±0.26 (3) 0.34 ±0.09 (5) 0.28 ±0.08 (4) 0.04 ±0.17 (1) 0.1 ±0.08 (0)
ROSCOE-GSM8K (2) 0.82 ±0.12 (2) 0.83 ±0.11 (2) 0.81 ±0.14 (2) 0.81 ±0.12 (2) 0.79 ±0.13 (2) 0.68 ±0.2 (2) 0.7 ±0.08 (2) 0.76 ±0.15 (2) 0.63 ±0.18 (2) 0.46 ±0.13 (2) 0.1 ±0.07 (1)
ROSCOE-eSNLI (2) 0.49 ±0.24 (2) 0.4 ±0.16 (2) 0.38 ±0.17 (2) 0.35 ±0.21 (2) 0.32 ±0.12 (2) 0.09 ±0.08 (0) 0.28 ±0.21 (1) 0.19 ±0.16 (1) 0.32 ±0.12 (2) 0.11 ±0.06 (0) 0.11 ±0.17 (1)
ROSCOE-DROP (2) 0.57 ±0.22 (2) 0.59 ±0.16 (2) 0.44 ±0.15 (2) 0.44 ±0.13 (2) 0.32 ±0.12 (2) 0.21 ±0.22 (1) 0.37 ±0.18 (2) 0.23 ±0.1 (2) 0.22 ±0.22 (1) 0.16 ±0.17 (1) 0.15 ±0.21 (1)
ROSCOE-CosmosQA (2) 0.57 ±0.18 (2) 0.55 ±0.18 (2) 0.51 ±0.16 (2) 0.57 ±0.17 (2) 0.53 ±0.21 (2) 0.33 ±0.25 (2) 0.48 ±0.17 (2) 0.44 ±0.26 (2) 0.57 ±0.2 (2) 0.13 ±0.04 (1) 0.49 ±0.24 (2)
NewsRoom (4) 0.59 ±0.02 (4) 0.59 ±0.03 (4) 0.44 ±0.05 (4) 0.55 ±0.03 (4) 0.5 ±0.07 (4) 0.36 ±0.06 (4) 0.16 ±0.05 (4) 0.45 ±0.04 (4) 0.26 ±0.06 (4) 0.21 ±0.08 (4) -0.01 ±0.04 (0)
SummEval (4) 0.35 ±0.06 (4) 0.44 ±0.14 (4) 0.54 ±0.08 (4) 0.38 ±0.02 (4) 0.48 ±0.02 (4) 0.19 ±0.06 (4) 0.13 ±0.06 (4) 0.29 ±0.09 (4) 0.4 ±0.12 (4) 0.15 ±0.05 (4) 0.06 ±0.02 (2)
WMT 2020 En-De (1) 0.63 (1) 0.37 (1) 0.51 (1) 0.46 (1) 0.2 (1) 0.42 (1) 0.15 (1) 0.11 (1) 0.36 (1) 0.15 (1) -0.03 (1)
WMT 2020 Zh-En (1) 0.54 (1) 0.39 (1) 0.48 (1) 0.41 (1) 0.25 (1) 0.42 (1) 0.15 (1) 0.14 (1) 0.39 (1) 0.15 (1) 0.01 (0)
WMT 2023 En-De (1) 0.22 (1) 0.14 (1) 0.23 (1) 0.16 (1) 0.17 (1) 0.22 (1) 0.19 (1) 0.08 (1) 0.18 (1) -0.09 (1) -0.05 (1)
WMT 2023 Zh-En (1) 0.17 (1) 0.14 (1) 0.19 (1) 0.14 (1) 0.15 (1) 0.15 (1) 0.14 (1) 0.02 (1) 0.15 (1) 0.01 (0) 0.01 (0)
Table 6: Scores per dataset for all models we evaluate: Cohen’s kappa for categorical annotations and Spearman’s correlation for graded annotations. For Spearman’s correlations, we report the number of significant correlations ($p<0.05$) for each model and dataset in brackets. Datasets in blue concern human-generated language while those in red concern model-generated text.
@@
Graded Annotations Dailydialog (1) 0.69 (1) 0.62 (1) 0.56 (1) 0.25 (1) 0.42 (1) 0.51 (1) 0.4 (1) 0.34 (1) 0.27 (1)
Switchboard (1) 0.6 (1) 0.48 (1) 0.53 (1) 0.07 (0) 0.38 (1) 0.17 (0) 0.36 (1) 0.35 (1) 0.08 (0)
Persona Chat (4) 0.2 ±0.09 (2) 0.09 ±0.2 (1) 0.11 ±0.06 (0) 0.06 ±0.15 (0) 0.17 ±0.21 (1) -0.04 ±0.22 (1) 0.04 ±0.13 (0) -0.01 ±0.22 (1) 0.06 ±0.18 (0)
Topical Chat (4) 0.22 ±0.02 (0) 0.14 ±0.13 (1) 0.11 ±0.1 (0) 0.06 ±0.18 (1) 0.1 ±0.14 (0) 0.16 ±0.17 (1) 0.25 ±0.05 (2) 0.17 ±0.09 (1) 0.08 ±0.13 (0)
Recipe-generation (6) 0.67 ±0.12 (6) 0.64 ±0.14 (6) 0.65 ±0.09 (6) 0.42 ±0.18 (5) 0.14 ±0.15 (2) 0.31 ±0.15 (2) 0.41 ±0.07 (6) 0.34 ±0.2 (4) 0.09 ±0.1 (0)
ROSCOE-GSM8K (2) 0.82 ±0.12 (2) 0.81 ±0.12 (2) 0.81 ±0.11 (2) 0.81 ±0.13 (2) 0.49 ±0.13 (2) 0.8 ±0.11 (2) 0.58 ±0.13 (2) 0.64 ±0.15 (2) 0.07 ±0.07 (0)
ROSCOE-eSNLI (2) 0.49 ±0.29 (2) 0.39 ±0.38 (1) 0.31 ±0.32 (1) 0.31 ±0.09 (2) 0.33 ±0.01 (2) 0.23 ±0.17 (1) 0.17 ±0.03 (1) 0.19 ±0.13 (1) -0.05 ±0.07 (0)
ROSCOE-DROP (2) 0.54 ±0.17 (2) 0.55 ±0.19 (2) 0.44 ±0.12 (2) 0.29 ±0.14 (2) 0.29 ±0.24 (1) 0.44 ±0.07 (2) 0.28 ±0.03 (2) 0.17 ±0.15 (1) -0.04 ±0.02 (0)
ROSCOE-CosmosQA (2) 0.57 ±0.21 (2) 0.55 ±0.22 (2) 0.56 ±0.11 (2) 0.55 ±0.12 (2) 0.62 ±0.15 (2) 0.38 ±0.06 (2) 0.54 ±0.21 (2) 0.32 ±0.14 (2) 0.3 ±0.14 (1)
NewsRoom (4) 0.57 ±0.05 (4) 0.53 ±0.03 (4) 0.53 ±0.05 (4) 0.46 ±0.02 (4) 0.19 ±0.06 (4) 0.49 ±0.04 (4) 0.19 ±0.04 (4) 0.22 ±0.12 (3) 0.09 ±0.06 (1)
SummEval (4) 0.45 ±0.11 (4) 0.48 ±0.16 (4) 0.33 ±0.03 (4) 0.35 ±0.05 (4) 0.17 ±0.05 (4) 0.24 ±0.13 (4) 0.38 ±0.1 (4) 0.24 ±0.11 (4) 0.03 ±0.07 (1)
WMT 2020 En-De (1) 0.57 0.44 (1) 0.38 (1) 0.39 (1) 0.13 (1) 0.34 (1) 0.33 (1) 0.3 (1) 0.0 (0)
WMT 2020 Zh-En (1) 0.52 (1) 0.44 (1) 0.44 (1) 0.42 (1) 0.19 (1) 0.36 (1) 0.39 (1) 0.35 (1) 0.13 (1)
WMT 2023 En-De (1) 0.2 (1) 0.16 (1) 0.18 (1) 0.2 (1) 0.2 (1) 0.18 (1) 0.21 (1) 0.19 (1) -0.01 (0)
WMT 2023 Zh-En (1) 0.15 (1) 0.13 (1) 0.15 (1) 0.16 (1) 0.16 (1) 0.13 (1) 0.11 (1) 0.13 (1) 0.06 (0)
Table 7: Scores per dataset for all models we evaluate using CoT prompts: Cohen’s kappa for categorical annotations and Spearman’s correlation for graded annotations. For Spearman’s correlation, we report the number of significant correlations for each model and dataset in brackets. Datasets in blue concern human-generated language while those in red concern model-generated text.